the distance between 2 stations is 300 km.Two motorcyclists start simultaneously from these stations and move towards each other.The speed of one of them is 7 km/h more than that of the other.If the distance between them after 2 hours of their start is 34 km, find the speed of each motorcyclists.

Answer by stanbon(75887)

the distance between 2 stations is 300 km.Two motorcyclists start simultaneously from these stations and move towards each other.The speed of one of them is 7 km/h more than that of the other.If the distance between them after 2 hours of their start is 34 km, find the speed of each motorcyclists.
-------------------
Faster-cyclist DATA:
rate = x + 7 km/h ; time = 2 hrs ; distance = 2(x+7) km
-----
Slower-cyclist DATA:
rate = x km/h ; time = 2 hrs ; distance = 2x km
----
Equation:
2(x+7) + 2x + 34 = 300
2x+14 + 2x = 66
4x = 52
x = 13 km/h speed of the slower
x + 7 = 20 km/h speed of the faster
======================================
Cheers,
Stan H.
============